# Fast-MCS: A Scalable Open-Source Tool to Find Minimal Cut Sets

### v1

@@ V Conclusions and Outlook
MCS (MCS) are crucial tools in evaluating a network’s resilience. In this work, we present our scalable, open-source tool, Fast-MCS, for efficiently obtaining MCS. We compared Fast-MCS’s performance against other methods, such as the combinatorial method and Shannon expansion, and show that Fast-MCS is faster than the other methods. Though the examples considered in this manuscript are communication network topologies, Fast-MCS is applicable to any system that can be represented as a graph. Currently, components are considered binary-valued, i.e., either working or unavailable. In the future, we intend to evaluate MCS as a time-varying function of components’ availabilities.

### v2

@@ V Conclusions and Outlook
Minimal Cut Sets (MCSs) are crucial tools in evaluating a network’s resilience. In this work, we present our scalable, open-source tool, Fast-MCS, for efficiently obtaining MCSs. We compared Fast-MCS’s performance against other methods, such as the combinatorial method and Shannon expansion, and show that Fast-MCS is faster than the other methods. Though the examples considered in this manuscript are communication network topologies, Fast-MCS is applicable to any system that can be represented as a graph. Currently, components are considered binary-valued, i.e., either working or unavailable. In the future, we intend to evaluate MCSs as a time-varying function of components’ availabilities.